Answer the following conceptual questions and practice about potential and kinetic energy.


1. You have two balls, a tennis ball and a baseball with twice the mass of the tennis ball. If you want to throw both balls with the same kinetic energy, how much faster should the tennis ball be compared to the baseball? Label your answer




1
Expert's answer
2021-11-23T10:32:13-0500

The kinetic energy of the baseball:


"KE=\frac12(2m)v^2=mv^2."

The kinetic energy of the tennis ball:


"KE=\frac12mu^2=mv^2,\\\space\\
u=v\sqrt2."